The body of mass 2kg undergoes a constant horizontal acceleration of 5m/s². Calculate the resultant horizontal force acting on the body. What will be the resultant force when it moves with a uniform velocity of 10m/s²
1
Expert's answer
2020-09-30T12:26:09-0400

According to the Newton's second law, the resultant foce acting on a body is equal to:


"F = ma"

where "m = 2kg" is the mass of the body and "a = 5m\/s^2" is its acceleration. Thus:


"F = 2\cdot 5 \space (kg\cdot m)/s^2 = 10N"

When abody moves with a uniform velocity, it does not have any acceleration ("a = 0" ). Thus, by the Newton's second law, the resultant force also is zero.


Answer. 1) 10N. 2) 0N.